parallel equation of y=x+9 (6,-2)
Answer by erica65404(394)

The slope of y = x + 9 is 1.
parallel lines have the same slope.
so the line you are looking for has a slope of 1.
y = x + b
to find b plug in the point they have given you.
-2 = 6 + b
-8 = b
Your final answer is y = x - 8